OpenStudy (anonymous):

How to rewrite r Cos(a t)+ s Sin(at) into single sinusoidal ? @M How to rewrite r Cos(a t)+ s Sin(at) into single sinusoidal ? @M @Mathematics

OpenStudy (anonymous):

hm i think i know it without the "a" part because i just looked this up the other week

OpenStudy (anonymous):

i believe its:\[A\sin(at+\phi)\]where \[A=\sqrt{r^2+s^2},\phi=\tan^{-1}\frac{s}{r}\]

OpenStudy (anonymous):

i might be wrong though, one sec.

OpenStudy (anonymous):

write it as \[\sqrt{r^2+s^2}\cos(t-\theta)\] where \[\tan(\theta)=\frac{s}{r}\]

OpenStudy (anonymous):

what joe said

OpenStudy (anonymous):

here is a reasonable explanation, but i saw a better one in a wiki that i cannot seem to find again http://www.mathcentre.ac.uk/resources/workbooks/mathcentre/web-rcostheta-alphaetc.pdf